OpenStudy (confusedstudenthere):

Solve for x. Enter your answer in interval notation using grouping symbols. x2 + 5x < 24

jimthompson5910 (jim_thompson5910):

First get everything to the same side. So subtract 24 from both sides \[\Large x^2 + 5x < 24\] \[\Large x^2 + 5x-24 < 24-24\] \[\Large x^2 + 5x-24 < 0\]

jimthompson5910 (jim_thompson5910):

what does \(\Large x^2 + 5x-24 \) factor to?

OpenStudy (confusedstudenthere):

I honestly have no idea. I'm so lost in this lesson

OpenStudy (confusedstudenthere):

(x+8)(x-3) ??

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

if you solved (x+8)(x-3) = 0, what two solutions do you get?

OpenStudy (confusedstudenthere):

(8,3) ?

jimthompson5910 (jim_thompson5910):

x = -8 is the solution to x+8 = 0 x = 3 is the solution to x-3 = 0 agreed?

OpenStudy (confusedstudenthere):

yes?

jimthompson5910 (jim_thompson5910):

(x+8)(x-3) = 0 means x+8 = 0 or x-3 = 0 solve each equation to get the two solutions x = -8 or x = 3

jimthompson5910 (jim_thompson5910):

so what we do is draw out a number line (see attached)

jimthompson5910 (jim_thompson5910):

Then we plot the two solutions (-8 and 3) on the number line See attached

jimthompson5910 (jim_thompson5910):

The red dots break up the number line into three regions region A in blue is from -infinity to -8 (excluding -8) region B is in green from -8 to +3 (excluding -8, excluding +3) region C is in purple from +3 to infinity (excluding +3) see attached

jimthompson5910 (jim_thompson5910):

with me so far @ConfusedStudentHere ?

OpenStudy (confusedstudenthere):

The third diagram confused me a bit as nothing in my lesson shows that type of graph. it only shows a green colored line like where the -8 to 3 is on those graphs you posted

jimthompson5910 (jim_thompson5910):

can you show me what you mean? please post a screenshot of how your lesson is showing it

OpenStudy (confusedstudenthere):

I can not as I have already moved on from this question but I still want to understand it.

jimthompson5910 (jim_thompson5910):

well the third diagram is just me color coding the 3 different regions so you can tell them apart

jimthompson5910 (jim_thompson5910):

anyways, can you tell me one number lies in the blue region A?

OpenStudy (zzr0ck3r):

another way to think about it (x+8)(x-3)<0 if and only if one of the two factors is negative and the other is positive case 1 x+8>0 and x-3<0 which gives x>-8 and x<3 so -8<x<3 case 2 x+8<0 and x-3>0 which gives x<-8 and x>3 and this cant happen. so the answer is -8<x<3